OpenStudy (anonymous):

Find the trigonometric form for (√3 + i)(1 + i).

OpenStudy (anonymous):

is this your question.. \[(\sqrt{3}+i)(1+i)\]

OpenStudy (anonymous):

no he's asking for the cis form

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

I know what he's asking..

OpenStudy (anonymous):

\[(\sqrt{3}+i)(1+i)=\sqrt{3}-1+(\sqrt{3}+1)i\]

OpenStudy (anonymous):

im trying to get it to a squared + b squared form , so i could put it in trigonometric form .

OpenStudy (anonymous):

I didnot finish yet..

OpenStudy (anonymous):

\[|z|=r=\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

\[x=\sqrt{3}-1\quad y=\sqrt{3}+1\]

OpenStudy (anonymous):

\[x^2=3-2\sqrt{3}+1\quad y^2=3+2\sqrt{3}+1\]

OpenStudy (anonymous):

\[x^2+y^2=8\] \[|z|=r=\sqrt{8}=2\sqrt{2}\]

OpenStudy (anonymous):

\[\theta=\tan^{-1} \frac yx\]

OpenStudy (anonymous):

\[\frac yx=2+\sqrt{3}\] \[\theta=\tan^{-1}( 2+\sqrt{3})=75\]

OpenStudy (anonymous):

\[z=rcis \theta\]

OpenStudy (anonymous):

\[z=2\sqrt{2}cis75\]

OpenStudy (anonymous):

or \[z=2\sqrt{2}cis \frac {\pi}{4}\]